NASA Space Station On-Orbit Status 9 May 2017 - Expedition 51 Crews Preps for Avionics Box Replacement Spacewalk

©NASA

NASA Space Station On-Orbit Status 9 May 2017 - Expedition 51 Crews Preps for Avionics Box Replacement Spacewalk.

The Expedition 51 crew is getting ready for a spacewalk Friday and working on several scientific investigations. Robotics controllers also swapped out a power relay box over the weekend through a complex and innovative robotic procedure.

Astronauts Peggy Whitson and Jack Fischer will exit the International Space Station for a 6.5-hour spacewalk on Friday. The duo will replace an avionics box that routes command and data information to external science experiments. Friday's spacewalk will be the 200th at the station for assembly and maintenance, the ninth for Whitson and the first for Fischer.

Whitson continued more research today comparing how bones adapt to space versus on Earth. Fischer stowed leaves that were harvested for the Veg-03 botany study and stowed them in a science freezer.

Friday evening and into Saturday, Robotics Ground Controllers maneuvered the Special Purpose Dexterous Manipulator (SPDM) Arms and successfully removed a failed Main Bus Switching Unit-2 and replaced it with a spare. The MBSU in question had stopped communicating telemetry back on April 25 but was still routing power to station systems.

The crew had installed a series of jumpers to power systems connected to the MBSU during the replacement, ensuring no impact to continued station operations. This was the first time an MBSU was swapped out by means other than a spacewalk. Since the successful replacement, the MBSU was powered up and checked out successfully with all station systems back to nominal power configuration.

On-Orbit Status Report

OsteoOmics: The crew changed BioCell media in BioCell Habitat 2. The media in the other four habitats will be changed this week. This is the second of four weeks of OsteoOmics operations. Crewmembers experience bone loss in orbit, stemming from the lack of gravity acting on their bones. OsteoOmics investigates the molecular mechanisms that dictate this bone loss by examining osteoblasts, which form bone, and osteoclasts, which dissolve bone. Improved understanding of these mechanisms could lead to more effective countermeasures to prevent bone loss during space missions and in a wide range of disorders on Earth. This may lead to better preventative care or therapeutic treatments for people suffering bone loss as a result of bone diseases like osteopenia and osteoporosis, or for patients on prolonged bed rest.

Veg-03: The crew harvested leaves from Chinese Cabbage growing in the Veggie facility and inserted samples into Minus Eighty Degree Celsius Laboratory Freezer for ISS (MELFI). The remainder of the plants were left to grow. The goal of Veg-03 is to further demonstrate the proof of concept for the Veggie plant growth chamber and planting pillows. Future long-duration missions into the solar system will require a fresh food supply to supplement crew diets, which entails growing crops in space. Previous investigations focused on improving productivity in controlled environments but the limited quarters of the space shuttle and ISS made it difficult to conduct large-scale crop production tests. Veg-03 expands on previous validation tests of the Veggie hardware to grow cabbage, lettuce and other fresh vegetables. Tests determine which types of microorganisms are present in space-grown cabbage, providing baseline data for future crop-growing efforts. Behavioral health surveys assess the impact of growing plants on crew morale and mood.

Extravehicular Activity (EVA) Preparations: The crew completed the following in preparation for the EXPRESS Pallet Controller Assembly (ExPCA) EVA currently planned for this Friday, May 12.

Prepared the Equipment Lock, Extravehicular Mobility Units (EMUs) and ancillary hardware to support suit donning.
Installed/replaced Metal Oxide (METOX) Lithium Hydroxide (LiOH) canisters and batteries in EMU 3006 and 3008.
Performed EMU pressurized fit verification to ensure proper fit and feel prior to the EVA.
Performed a procedures review covering reminders, emergency briefing and helpful tips. Following the review, the crew completed a conference with ground teams to address any questions or concerns.
Orbital ATK (OA)-7 Cargo Operations: Yesterday afternoon, the crew continued transferring Cygnus cargo to ISS. Approximately 3 hours of transfer operations remain to be completed. Cygnus is currently scheduled to unberth from ISS on July 16th.

Today's Planned Activities
All activities were completed unless otherwise noted.

Personal Data Prep for Return
Atmospheric Control System (ACS) NORS Oxygen Teardown
Test Video Recording for Russia Today TV Channel
Preparation of Reports for Roscosmos Web Site and Social Media
URAGAN. Observation and photography
EKON-M. Observations and photography
Extravehicular Activity (EVA) Reminder for On-Orbit Fit check Verification (OFV)
SARCOLAB-3 CONSTRAINTS
Acoustic Dosimeter Setup Day 2
OsteoOmics Media Relocate
VEG-03 Science Harvest
OsteoOmics Media Change
VEG-03 MELFI Insertion #1
On-Orbit Hearing Assessment (O-OHA) with EarQ Software Setup and Test
Nanoracks Cubesat Deployer PAO Video
Extravehicular Activity (EVA) Equipment Lock Preparation Part 1
Extravehicular Mobility Unit (EMU) Metal Oxide (METOX) Installation
Extravehicular Mobility Unit (EMU) On-orbit Fit check Verification
Приватная семейная конференция
CASKAD. Manual Mixing in Bioreactor
Extravehicular Activity (EVA) Procedure Review
Extravehicular Activity (EVA) Procedure Conference
Water Recovery System Waste Water Tank Drain Init
On-Orbit Hearing Assessment (O-OHA) with EarQ Software Setup and Test
Water Recovery System Waste Water Tank Drain Termination
Early Mission Debrief Conference
OsteoOmics MELFI Removal

Completed Task List Items
Veggie 03 Pillow Watering and Photo

Ground Activities
All activities were completed unless otherwise noted.
EVA prep support
SPDM Stow and MT Translation from WS3 to WS4
MSS ExPCA EVA Setup

Three-Day Look Ahead:
Wednesday, 05/10: EVA procedures review, MARES, OsteoOmics media change, OBT ISS emergency simulation
Thursday, 05/11: EVA Equipment Lock prep part 2, EVA procedures conference, EVA tool audit, OsteoOmics,
Friday, 05/12: ExPCA EVA #42

QUICK ISS Status - Environmental Control Group:
Component - Status
Elektron - On
Vozdukh - Manual
[СКВ] 1 - SM Air Conditioner System ("SKV1") - Off
[СКВ] 2 - SM Air Conditioner System ("SKV2") - On
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Operate
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) - Process
Urine Processing Assembly (UPA) - Process
Trace Contaminant Control System (TCCS) Lab - Full up
Trace Contaminant Control System (TCCS) Node 3 - Off

Please follow SpaceRef on Twitter and Like us on Facebook.